OpenStudy (anonymous):

3. A rectangular pasture has a fence around the perimeter. The length of the fence is 16x7 and the width is 48x4. What is the area of the pasture?

OpenStudy (anonymous):

So, the length is 16 * 7 and width is 48 *4?

OpenStudy (anonymous):

I think this is \[768x ^{11}\]

OpenStudy (anonymous):

And yes it is. I just worked the problem. Not sure if I did it right or not.

OpenStudy (anonymous):

Well, just multiply those two numbers together. So (16*7) * (48*4) = l*w = 112*192=21504 is area in w/e unit. so ft^2, m^2, cm^2, yd ^2)

OpenStudy (anonymous):

Oh im so confused.

OpenStudy (anonymous):

so my answer wasnt right?

OpenStudy (anonymous):

Is it an x? or a multiply sign???

OpenStudy (anonymous):

its an x not a multiply sign

OpenStudy (anonymous):

Okay then, it's different, please use the equation editor okay?\[ 16x^7 * 48x^4 = 768x^{11}\]

OpenStudy (anonymous):

okay this is my whole problem. \[The length of the fence is 16x ^{7} and the width is 48x ^{4}\]

OpenStudy (anonymous):

I just solved it above. Anyhow, next time. Just use normal text. But when you need ONLY the math part, use the equation editor.

OpenStudy (anonymous):

okay i will. heres another one. 4. What is the simplified form of written in scientific notation? \[(3.25 × 10^{3})(7.8 × 10^{6})\]

OpenStudy (anonymous):

Just expand it out: You can do this on any calc. Even google. 25350000000

OpenStudy (anonymous):

what do you mean? these are my options. \[2.535 × 10^{10}\] \[2.535 × 10^{18}\] \[2.535 × 10^{19}\] \[25.35 × 10^{9}\]

OpenStudy (anonymous):

Oh, just convert it to the appropriate scientific notation then, which is: option 1.

OpenStudy (anonymous):

can you tell me how you got option 1? cause I got option 2. because I multipled the exponents.

OpenStudy (anonymous):

You can't do that for this. You have to expand them out, multiply those numbers, and then put the answer in scientific notation.

OpenStudy (anonymous):

oh goodness, I did not know that. But I do now. Okay now all I have is one more problem and I'll be done. Can you help me with it??

OpenStudy (anonymous):

nvm. I got the answer :)